OpenStudy (anonymous):

The Slope of the line normal to the curve e^x- x^3+ y^2= 10 at the point (0, 3) is

OpenStudy (asnaseer):

First use implicit differentiation to calculate an expression for the slope at any point (x,y)

OpenStudy (anonymous):

@asnaseer I got (y'=-e^x+ 3x^2)/2y

OpenStudy (asnaseer):

that is correct. Now use this to calculate the slope to the curve at the point (0, 3) Finally use the fact that the product of the slopes of to lines that are perpendicular to each other is -1 to work out the slope of the normal

OpenStudy (anonymous):

so @asnaseer would I plug in (0, 3)... (y' -e^0 + 3*0^2)/ 2*3 ?

OpenStudy (asnaseer):

yes, assuming you meant y'=(-e^0+3*0^2)/2*3

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

then you're good to go :)

OpenStudy (anonymous):

@asnaseer the answer seems to be -1/6 did I do it correctly?

OpenStudy (anonymous):

or -x/6

OpenStudy (asnaseer):

yes - that is the slope to the curve at (0, 3) now you just need to calculate the slope of the normal line using the technique I mentioned above :)

OpenStudy (asnaseer):

how can it be -x/6?

OpenStudy (anonymous):

I don't know that's what the calculator said but I change it to -1/6

OpenStudy (asnaseer):

there is no 'x' in the expression (-e^0+3*0^2)/2*3

OpenStudy (asnaseer):

and, you don't need a calculator for this

OpenStudy (anonymous):

I do, How do I do what you said above?

OpenStudy (asnaseer):

take each term: what is e^0=?

OpenStudy (anonymous):

1?

OpenStudy (asnaseer):

correct (because any number to the power of zero is one) next what is: 3*0^2 = ?

OpenStudy (anonymous):

0

OpenStudy (asnaseer):

correct (because anything times zero is zero) finally, what is: 2*3 = ?

OpenStudy (anonymous):

8

OpenStudy (asnaseer):

are you sure?

OpenStudy (anonymous):

Oh sorry I read it as 2^3 that would be 6

OpenStudy (asnaseer):

good, so we end up with: y' = (-e^0+3*0^2)/2*3 = (-1 + 0)/6 = -1/6 no calculator required :)

OpenStudy (anonymous):

Oh yeah your right :) guess I'm getting to use to a calculator. There is another part to do?

OpenStudy (asnaseer):

this just gives you the slope to the curve at the point (0,3) you are asked to find the slope of the NORMAL - i.e. the slope of the line that would be perpendicular to this line

OpenStudy (anonymous):

okay so in my head I would take it back to algebra and tell myself to find the negative reciprocal, but this is calc so I have to do something else don't I?

OpenStudy (asnaseer):

no - don't try to complicate things just because you are in calc :) your 1st instinct is correct :)

OpenStudy (anonymous):

Oh gees easiest thing I've done all day god I miss algebra. So the final answer would be 6

OpenStudy (asnaseer):

the "calc" part was used to calculate the initial expression for the slope at any point

OpenStudy (asnaseer):

perfect! :)

OpenStudy (anonymous):

Thank you one billion times

OpenStudy (asnaseer):

you are more than welcome my friend :)